# Entanglement Dynamics in Monitored Systems and the Role of Quantum Jumps

### v2

@@ Abstract
Monitored quantum many-body systems display a rich pattern of entanglement dynamics, which is unique to this non-unitary setting. This work studies the effect of quantum jumps on the entanglement dynamics beyond the no-click limit corresponding to a deterministic non-Hermitian evolution. We consider two examples, a monitored SSH model and a quantum Ising chain, for which we show the jumps have remarkably different effects on the entanglement despite having the same statistics as encoded in their waiting-time distribution. To understand this difference, we introduce a new metric, the statistics of entanglement gain and loss due to jumps and non-Hermitian evolution. This insight allows us to build a simple stochastic model of a random walk with partial resetting, which reproduces the entanglement dynamics, and to dissect the mutual role of jumps and non-Hermitian evolution on the entanglement scaling. We demonstrate that significant deviations from the no-click limit arise whenever quantum jumps strongly renormalize the non-Hermitian dynamics, as in the case of the SSH model at weak monitoring or in the Ising chain at large transverse field. On the other hand, we show that the weak monitoring phase of the Ising chain leads to a robust sub-volume logarithmic phase due to weakly renormalized non-Hermitian dynamics.
@@ I Introduction
Entanglement transitions due to measurements have been studied broadly in two somewhat different settings. On the one hand, the stochastic dynamic encoded in a quantum many-body trajectory describes the system’s evolution conditioned to a set of measurement outcomes. A crucial aspect of this criticality is that it is hidden in the rare fluctuations of the measurement process, probed by non-linear functional of the state such as the entanglement entropy or the purity in a dynamical purification protocol Gullans and Huse (2020), while conventional observables averaged over the noise are usually transparent to it. This makes the theoretical description and experimental detection of MIPT particularly challenging, even though recent progress has been made Noel et al. (2022); Koh et al. (2023); Google AI and Collaborators (2023). On the theoretical front volume-to-area law entanglement transitions have been reported in monitored random circuits Li et al. (2018, 2019); Skinner et al. (2019); Szyniszewski et al. (2019); Jian et al. (2020); Choi et al. (2020); Zabalo et al. (2022); Sierant and Turkeshi (2022); Sierant et al. (2022); Klocke and Buchhold (2023) and non-integrable Hamiltonian Fuji and Ashida (2020); Lunt and Pal (2020); Doggen et al. (2022); Xing et al. (2023); Altland et al. (2022) with projective or weak measurements. Monitored non-interacting systems, on the other hand, are not expected to sustain a volume law phase Cao et al. (2019); Fidkowski et al. (2021). Still, a critical sub-volume phase, whose origin and stability are currently under debate Coppola et al. (2022); Lóio et al. (2023); Poboiko et al. (2023a); Jian et al. (2023); Fava et al. (2023); Carisch et al. (2023); Jin and Martin (2023), has been numerically found in several works Alberton et al. (2021); Van Regemortel et al. (2021); Turkeshi et al. (2021); Botzung et al. (2021); Bao et al. (2021); Turkeshi et al. (2022a); Piccitto et al. (2022); Kells et al. (2023); Paviglianiti and Silva (2023); Müller et al. (2022); Buchhold et al. (2022).
@@
The relation between these two limits of the measurement problem, particularly concerning the entanglement dynamics, is not well understood. Is the no-click dynamics stable enough to include QJs, which should be seen as irrelevant perturbations? Or, on the contrary, do QJs completely change the entanglement structure of the monitored system? These questions ultimately go beyond measurement-induced transitions and touch upon the relevance of non-Hermitian Hamiltonians in the description of open quantum systems Minganti et al. (2019); Gneiting et al. (2021); Naghiloo et al. (2019); Han et al. (2023).
@@
This work studies the quantum jumps dynamics of two monitored free-fermion models, the Su-Schrieffer-Heeger (SSH) model and the Quantum Ising chain. These models are attractive for two reasons: first, they differ in their global symmetry, a fact which is believed to play a critical role in the stability of the MIPT; second, their entanglement transition in the no-click limit has been studied in detail Turkeshi and Schiró (2023); Gal et al. (2023). By solving the full stochastic dynamics and computing the entanglement entropy, we show that QJs have remarkably different impacts on their respective phase diagram as compared to the no-click limit. This is particularly surprising as the two models share the same jump statistics, as we show by computing the QJs waiting-time distribution. To track down the origin of this difference we introduce a new metric that captures the effect of QJs on the entanglement. Specifically we compute the statistics of entanglement entropy changes after and in-between quantum jumps. We show that, on average, QJs induce an entanglement entropy loss while the non-Hermitian evolution causes a gain, although the statistics display rather broad tails and fluctuations.
Using this insight we introduce a classical stochastic random walk model with partial resetting, built-in using the full statistics of entanglement gain and loss, which is able to reproduce the full QJ dynamics of entanglement entropy for both models. More importantly, the entanglement gain and loss picture clarifies the mutual role of quantum jumps and non-Hermitian evolution in the scaling of entanglement entropy. Our analysis reveals that a crucial role of quantum jumps is to renormalize the effective non-Hermitian dynamics with respect to the bare no-click limit. We show that this renormalization, or its lack thereof, can naturally explain the different behaviors of quantum jumps in the two models and provide a criterion for the relevance of the no-click limit. Our results suggest a form of universality in how quantum jumps affect entanglement, providing a framework for interpreting the stochastic many-body dynamics in Hilbert space, which opens up to several applications.
@@
The paper is structured as follows. In Sec. II we present a summary of the main results of this work. In Sec. III we review the background material: we introduce the QJ measurement protocol and the models we will considered throughout this work as well as the quantities we will use to characterize them. In Sec. IV present our numerical results for the entanglement dynamics of the monitored SSH and Ising chain. In Sec. V we present our results for the waiting time distribution of quantum jumps. In Sec. VI we introduce our new metric, the statistics of entanglement gain and loss and describe a classical stochastic model for the entanglement dynamics. In Sec. VII we use the insights from this model to discuss the role of jumps and non-Hermitian dynamics on the scaling of the entanglement and the MIPT. Finally Sec. VIII contains our conclusions and future perspectives. In the Appendixes we provide further methodological details and results relevant for our work.
Figure 1: Panel (a) cartoon of the setup. A monitored fermionic chain evolving under quantum jump protocol. Panel (b): A quantum trajectory we obtain from this type of setup. The dots symbolize the jumps, and the inset shows how the quantities $\Delta S_{\rm qj}$ (resp. $\Delta S_{\rm nh}$) are extracted from the change of entanglement happening in a jump (resp. during the non-Hermitian evolution). We define the effective slope $\delta_{\mathrm{nH}}$ of the entanglement due to the non-Hermitian evolution between two consecutive jumps. The local entanglement $S$ at which these events happen is stored to plot the joint probability of the following panels (c-d). In panel (c) is plotted in plain line the distribution of $\Delta S_{qj}$, and the dots represent the joint version of this distribution with respect to the entanglement when the jumps is happening. In panel (d) we use the same typology but plotting this time the quantity $\delta S_{\mathrm{nH}}$ which is the effective non-Hermitian slope defined before. Panel (e): Cartoon depicting the entanglement dynamics as a classical random walk (RW) with stochastic drift and resetting drawn from the histograms in panels (c-d). Panel (f): Comparison between the dynamics of the average entanglement entropy obtained from the exact quantum jump dynamics (QJ) or the phenomenological classical random walk (RW).
II Overview
In this Section, we summarize the main results of this work. We consider the entanglement dynamics of two models of non-interacting monitored fermions (See Fig. 1 (a) for a sketch), the Su-Schrieffer-Heeger (SSH) chain and the Quantum Ising chain, in the presence of Quantum Jump (QJ) monitoring. These models, in their no-click limit, feature a rich phase diagram with entanglement transitions separating a volume (SSH)/sub-volume (Ising) entanglement phase at weak monitoring from an area law phase at strong monitoring Turkeshi and Schiró (2023); Gal et al. (2023). In Sec. IV, we compute the entanglement entropy under full stochastic monitoring dynamics and show that the effect of QJs is remarkably different for the two models. Specifically, we show that jumps in the SSH model have a substantial impact and qualitatively change the phase diagram with respect to the no-click limit. This fact is particularly manifest at weak monitoring, where the volume law phase is turned into a much less entangled, possibly logarithmic, phase. Such a phase eventually turns into an area law phase above a critical monitoring strength, whose value is also strongly renormalized by the jumps. In the Ising chain, on the other hand, our results show that QJs modify mainly quantitatively the no-click phase diagram and overall appear to be less of a relevant perturbation, except at large values of the transverse field, even though compelling differences emerge between the two phase diagrams.
In Sec. V, to understand the origin of such a different impact of QJs on the dynamics of the two models, we study the statistics of quantum jumps, particularly their waiting time distribution. The statistics of waiting times, which is naturally accessible using a higher-order Montecarlo wave function scheme, gives information on the typical time scale between jumps, i.e., the time scale during which the non-Hermitian evolution is at play (See Fig. 1 (b)). Surprisingly, in both models, QJs at long times approach a Poisson distribution of waiting times with the same average waiting time. This analysis suggests that the difference observed in the entanglement dynamics does not arise from how jumps are distributed in time.
To proceed further, we introduce a new metric, namely the statistics of the changes to the entanglement entropy after and in between QJs. Specifically, we consider the stochastic dynamics of the entanglement entropy along a quantum trajectory (See Fig. 1 (b)) : the typical pattern is given by an evolution driven by the non-Hermitian Hamiltonian interrupted abruptly by a discontinuous change in entanglement entropy due to a quantum jump. The entanglement entropy can either increase or decrease after a jump and, similarly, can either grow or diminish during the time between the jumps, where non-Hermitian evolution occurs. Given this pattern, we now ask : what is the statistics of the entanglement entropy change after a quantum jump and after an evolution with the non-Hermitian Hamiltonian ? To answer this question, we sample the joint probability of observing a change $\Delta S_{\rm qj}$ due to the jumps or $\delta S_{\rm nH}$ due to the non-Hermitian evolution (see inset of Fig. 1 (b)) together with a particular value of entanglement entropy $S$, i.e., $P(\Delta S_{\rm qj},S),P(\delta S_{\rm nh},S)$ which are obtained by binning stochastic events (jumps or non-Hermitian evolution) according to the entanglement content of the state they act on (which means the entanglement of the state before the event happens). The resulting histograms, plotted in Fig. 1 (c-d) for the Ising chain, are strongly asymmetric confirming the intuition that QJs induce on average an entanglement loss, while the non-Hermitian dynamics an entanglement gain, as we discuss in Sec. VI. Furthermore they feature broad tails which depend strongly on the entanglement content of the system, indicating for example that highly entangled states are more fragile to QJs.
Using the full statistics of entanglement changes we propose in Sec. VI a classical stochastic model (See Fig. 1 (e)) to understand the entanglement dynamics as a random walk with drift and partial resetting Evans and Majumdar (2011); Evans et al. (2020); Turkeshi et al. (2022a)). Within this model the probability of having an entanglement entropy $S$ at time $t$, that we note $\mathcal{P}_{t}(S)$, evolves according to the master equation
$\begin{split}&\mathcal{P}_{t+dt}(S)=rdt\int_{\Delta S_{\rm qj}}P(\Delta S_{\rm qj},S-\Delta S_{\rm qj})\mathcal{P}_{t}(S-\Delta S_{\rm qj})+\\ &+(1-rdt)\int_{\delta S_{\rm nH}}P(\delta S_{\rm nH},S-\delta S_{\rm nH}dt)\mathcal{P}_{t}(S-\delta S_{\rm nH}dt)\end{split}$ (1)
where the first term describes the jump, which adds a random contribution $\Delta S_{\rm qj}$ with probability $rdt\times P(\Delta S_{\rm qj},S-\Delta S_{\rm qj})$, with $r$ the resetting rate, while the second one describes the non-Hermitian dynamics which increase the entanglement of a random slope $\delta S_{\rm nH}$ with probability $(1-rdt)\times P(\delta S_{\rm nH},S-\delta S_{\rm nH}dt)$. Using this classical stochastic model, we can reproduce accurately the dynamics of the average entanglement entropy as a function of the monitoring rate (See Fig. 1 (f) for the Ising chain).
In Sec. VII we use the insights obtained from our stochastic model to analyze the results for the entanglement in the two models and clarify the role of QJs and non-Hermitian evolution in the scaling of the steady-state entanglement entropy. Our analysis reveals, for example, that the logarithmic phase for the monitored SSH chain is mainly due to quantum jumps. At the same time, the non-Hermitian dynamics is strongly renormalized compared to the no-click limit. On the other hand, for the Ising chain at small values of monitoring rate and field, we see a logarithmic phase arising from both jumps and non-Hermitian evolution, the latter appearing to be only weakly renormalized with respect to the no-click limit. Upon increasing the transverse field or the monitoring rate, however, we find evidence of a strong renormalization of the non-Hermitian dynamics due to the jumps, which explains the observed departure from the no-click limit. We conclude our analysis in Sec. VII by discussing the implications of our findings for the MIPT of free fermions.
III Background
@@
In this Section, we summarize relevant background material, including the description of the Quantum Jump protocol and its sampling via the Montecarlo Wave Function and the Waiting Time-Distribution, and introduce the models we focus on throughout the paper and the methods to solve their dynamics. Finally, we recall their entanglement transition in the no-click limit and put that in the light of known results of monitored fermions.
III.1 Monitoring by Quantum Jumps
We start from the general Quantum Jump protocol for a given set of Kraus operators $\left\{M_{\alpha}\right\}$, for which $\sum_{\alpha\geq 0}M_{\alpha}M_{\alpha}^{\dagger}=\mathbb{1}$ Breuer and Petruccione (2002); Wiseman and Milburn (2009). The evolution preserves the purity and is given by
$|\psi(t+dt)\rangle=\sum_{\alpha\geq 0}d\xi_{\alpha}\frac{M_{\alpha}|\psi(t)\rangle}{\sqrt{\langle\psi(t)|M_{\alpha}^{\dagger}M_{\alpha}|\psi(t)\rangle}}$ (2)
where $d\xi_{0}\equiv(1-\sum_{\alpha>0}d\xi_{\alpha})$, while, for $\alpha>0$, $d\xi_{\alpha}\in\{0,1\}$ is an increment for the inhomogeneous Poisson process with average ${P(d\xi_{\alpha}=1)=\langle\psi(t)|M_{\alpha}^{\dagger}M_{\alpha}|\psi(t)\rangle}.$ At infinitesimal times $dt$, the Kraus operators are expressed in terms of the Hamiltonian $H$ and of jump operators $\{L_{\alpha}\}$ as $M_{0}=1-dt(iH+\frac{1}{2}\sum_{\alpha>0}L_{\alpha}^{\dagger}L_{\alpha})$ and $M_{\alpha}=\sqrt{dt}L_{\alpha}$. We note that the $\frac{1}{2}\sum_{\alpha>0}L_{\alpha}^{\dagger}L_{\alpha}$ in $M_{0}$ is imposed by the normalization condition. It is then possible to obtain a Stochastic Schrödinger Equation using Itô calculus ($d\xi_{\alpha}\propto\sqrt{dt}$, $d\xi_{\alpha}M_{0}=d\xi_{\alpha}+o(dt)$) and keeping first order terms in $dt$, which reads Dalibard et al. (1992); Plenio and Knight (1998); Daley (2014)
$d|\psi(t)\rangle$ $=-idt\left\{H-\frac{i}{2}\sum_{\alpha>0}(L_{\alpha}^{\dagger}L_{\alpha}-\langle L_{\alpha}^{\dagger}L_{\alpha}\rangle_{t})\right\}|\psi(t)\rangle$
$\ +\sum_{\alpha>0}d\xi_{\alpha}\left\{\frac{L_{\alpha}}{\sqrt{\langle L_{\alpha}^{\dagger}L_{\alpha}\rangle}}-1\right\}|\psi(t)\rangle.$ (3)
@@ III.1 Monitoring by Quantum Jumps
This last equation is composed of a deterministic part which corresponds to an effective non-Hermitian Hamiltonian $H_{\rm eff}$ given by
$H_{\mathrm{eff}}=H-\frac{i}{2}\sum_{\alpha>0}L_{\alpha}^{\dagger}L_{\alpha}$ (4)
@@
and a stochastic one due to the action of quantum jumps. The effective Hamiltonian is not Hermitian because of the back-action of the jumps and corresponds to the no-click limit, where no jumps happen during the trajectory. We note that the evolution of the system is state-dependent (thus nonlinear); see the counter-term appearing in Eq. (III.1), to ensure the normalization.
III.2 Monte Carlo Wave Function and Waiting Times
The stochastic Schrödinger equation (III.1) is solved via Monte Carlo methods Daley (2014), which we briefly review below starting from the first-order integration schemes. In this case, the idea is to divide the time evolution into segments of interval $dt$, and act on the system accordingly to the probability of having a quantum jump $P_{\mathrm{Jump}}\equiv\sum_{\alpha>0}P(d\xi_{\alpha}=1)$ or evolving under the non-Hermitian Hamiltonian $P_{\mathrm{nH}}\equiv 1-P_{\mathrm{Jump}}$. At a sufficiently small time interval $dt$, the probability of a jump $P_{\mathrm{Jump}}$ is small, since $P(d\xi_{\alpha}=1)=dt\langle\Psi(t)|L_{\alpha}^{\dagger}L_{\alpha}|\Psi(t)\rangle$ with $\alpha>0$. If a jump occurs, we divide the interval [0,1] into segments of length $P(d\xi_{\alpha}=1)/P_{\mathrm{jump}}$. The jump channel $\tilde{\alpha}>0$ is selected checking in which subinterval a uniformly drawn random number falls. Then, the measurement is assumed to occur in $dt$ time, and leads to
$|\Psi(t+dt)\rangle=\frac{L_{\tilde{\alpha}}|\Psi(t)\rangle}{\sqrt{\langle\Psi(t)|L_{\tilde{\alpha}}^{\dagger}L_{\tilde{\alpha}}|\Psi(t)\rangle}}.$ (5)
If no jump occurs, the state is given by
$|\Psi(t+dt)\rangle=\frac{(1-iH_{\mathrm{eff}}dt)|\Psi(t)\rangle}{\sqrt{P_{\mathrm{nH}}}}=\frac{|\tilde{\Psi}(t)\rangle}{\sqrt{P_{\mathrm{nH}}}}$ (6)
This process is the implementation of the POVM described before (cf. Eq. (2)) with a finite $dt$. We note that $P_{\mathrm{nH}}=1-P_{\mathrm{Jump}}$ is due to
$\begin{split}&\langle\tilde{\Psi}(t)|\tilde{\Psi}(t)\rangle=1-dt\left\langle\Psi(t)\left|\mathrm{i}\left(H_{\mathrm{eff}}-H_{\mathrm{eff}}^{\dagger}\right)\right|\Psi(t)\right\rangle\\ &\quad=1-dt\sum_{\alpha>0}\left\langle\Psi(t)\left|L_{\alpha}^{\dagger}L_{\alpha}\right|\Psi(t)\right\rangle=1-P_{\mathrm{Jump}}\end{split}$ (7)
which is a consequence of the normalisation of the Kraus operators in the POVM Breuer and Petruccione (2002); Wiseman and Milburn (2009). This procedure has the drawback of not having a natural way to control the accuracy of the simulation, which is instead empirically benchmarked by considering different choices of $dt$ and checking that there is not qualitative difference in disorder averages.
We overcome these limitation considering a higher-order scheme that is based on sampling the time at which subsequent jumps happen Breuer and Petruccione (2002); Daley (2014), i.e., using the cumulative waiting-time distribution defined as
@@ III.2 Monte Carlo Wave Function and Waiting Times
Specifically, we proceed by iterating the following loop : (i) Assume a normalised state $|\Psi(t_{i})\rangle$ is reached at time $t$ (possibly the initial time of the dynamics, in which case $|\Psi(t)\rangle$ is the initial state). (ii) Extract a random number $r$ uniformly distributed in $[0,1]$ and find the random waiting time $\tau$ for the next jump by solving the equation $r=1-F[\Psi(t),\tau]$.
@@
(iv) At time $t+\tau$ a quantum jump occurs. As before, the output channel is chosen splitting the [0,1] interval into segments of size $\langle\Psi(t+\tau)|L^{\dagger}_{\alpha}L_{\alpha}|\Psi(t+\tau)\rangle$ and checking in which one a uniformly drawn random number falls. The jump is immediate and the post-measurement state is
$|\Psi(t+\tau^{+})\rangle=\frac{L_{\tilde{\alpha}}|\Psi(t+\tau)\rangle}{\sqrt{\langle\Psi(t+\tau)|L_{\tilde{\alpha}}^{\dagger}L_{\tilde{\alpha}}|\Psi(t+\tau)\rangle}}.$ (10)
@@
In a nutshell, a quantum trajectory is specified as a sequence of non-Hermitian quantum quenches interspersed with discontinuous jumps, that set the initial conditions for the forecoming integration steps.
An advantage of this approach is to access directly the waiting-time distribution (WTD) or delay function Cohen-Tannoudji and Dalibard (1986); Daley (2014); Landi et al. (2023). We will discuss the behavior of the waiting time distribution for the models of interest later on. In general, computing the waiting time distribution is non-trivial and requires the knowledge of the stochastic process. There is, however, a special case where this can be done. Whenever the imaginary part of the non-Hermitian Hamiltonian commutes with the real part, then one can conclude that the norm decay is exponential and therefore so the cumulative waiting time-distribution, $F=1-\exp\left(-\Gamma\tau\right)$. In this case, the time for the next jump can be directly obtained as a function of a random number $r$, as $\tau=-\frac{1}{\Gamma}\mbox{ln}r$. Such an instance occurs for a system with local density monitoring and conservation of global particle number Alberton et al. (2021); Coppola et al. (2022).
III.3 Average vs Conditional Dynamics and Entanglement Entropy
In the following we will be interested in the conditional jump dynamics described by the stochastic Schrödinger equation (III.1) and in particular in the entanglement dynamics of the conditional state, also called a quantum trajectory. Indeed, it is known that by averaging the density matrix over the stochastic noise in Eq. (III.1) one recovers a Lindblad master equation, which describes in the present context heating towards a trivial infinite temperature state. In the following we will be mainly interested in the (von Neumann) entanglement entropy, defined as  Calabrese and Cardy (2004); Amico et al. (2008)
$S(\mathbf{\xi}_{t})=-\mathrm{tr}_{A}\left[\rho_{A}(\mathbf{\xi}_{t})\ln\rho_{A}(\mathbf{\xi}_{t})\right]\;,$ (11)

### v3

@@ Abstract
Monitored quantum many-body systems display a rich pattern of entanglement dynamics, which is unique to this non-unitary setting. This work studies the effect of quantum jumps on the entanglement dynamics beyond the no-click limit corresponding to a deterministic non-Hermitian evolution. To this aim we introduce a new tool which looks at the statistics of entanglement entropy gain and loss after and in-between quantum jumps. This insight allows us to build a simple stochastic model of a random walk with partial resetting, which reproduces the entanglement dynamics, and to dissect the mutual role of jumps and non-Hermitian evolution on the entanglement scaling. We apply these ideas to the study of measurement-induced transitions in monitored fermions. We demonstrate that significant deviations from the no-click limit arise whenever quantum jumps strongly renormalize the non-Hermitian dynamics, as in the case of models with $U(1)$ symmetry at weak monitoring. On the other hand, we show that the weak monitoring phase of the Ising chain leads to a robust sub-volume logarithmic phase due to weakly renormalized non-Hermitian dynamics.
@@ I Introduction
Entanglement transitions due to measurements have been studied broadly in two somewhat different settings. On the one hand, in the stochastic dynamics encoded in a quantum many-body trajectory describing the system’s evolution conditioned to a set of measurement outcomes. In this setting, the criticality is hidden in the rare fluctuations of the measurement process, probed by a non-linear functional of the state such as the entanglement entropy or the purity in a dynamical purification protocol Gullans and Huse (2020), while conventional observables averaged over the noise are usually transparent to it. This makes the theoretical description and experimental detection of MIPT particularly challenging, even though recent progress has been made Noel et al. (2022); Koh et al. (2023); Google AI and Collaborators (2023). On the theoretical front volume-to-area law entanglement transitions have been reported in monitored random circuits Li et al. (2018, 2019); Skinner et al. (2019); Szyniszewski et al. (2019); Jian et al. (2020); Choi et al. (2020); Zabalo et al. (2022); Sierant and Turkeshi (2022); Sierant et al. (2022); Klocke and Buchhold (2023) and non-integrable Hamiltonian Fuji and Ashida (2020); Lunt and Pal (2020); Doggen et al. (2022); Xing et al. (2023); Altland et al. (2022) with projective or weak measurements. Monitored non-interacting systems, on the other hand, are not expected to sustain a volume law phase Cao et al. (2019); Fidkowski et al. (2021). Still, a critical sub-volume phase, whose origin and stability are currently under debate Coppola et al. (2022); Lóio et al. (2023); Poboiko et al. (2023); Jian et al. (2023); Fava et al. (2023); Carisch et al. (2023); Jin and Martin (2023), has been numerically found in several works Alberton et al. (2021); Van Regemortel et al. (2021); Turkeshi et al. (2021); Botzung et al. (2021); Bao et al. (2021); Turkeshi et al. (2022); Piccitto et al. (2022); Kells et al. (2023); Paviglianiti and Silva (2023); Müller et al. (2022); Buchhold et al. (2022).
@@
The relation between these two limits of the measurement problem, particularly concerning the entanglement dynamics, is not well understood. Is the no-click dynamics stable enough to the inclusion of QJs, which should be seen as irrelevant perturbations? Or, on the contrary, do QJs completely change the entanglement structure of the monitored system? These questions ultimately go beyond measurement-induced transitions and touch upon the relevance of non-Hermitian Hamiltonians in the description of open quantum systems Yuto Ashida and Ueda (2020); Lau and Clerk (2018); Minganti et al. (2019); Naghiloo et al. (2019); Minganti et al. (2020); Han et al. (2023); Ehrhardt and Larson (2023).
To understand the role of QJs and non-Hermitian evolution on the entanglement dynamics in monitored systems we introduce a new tool: the statistics of the entanglement entropy changes after and in-between QJs. We show that, on average, QJs induce an entanglement entropy loss while the non-Hermitian evolution causes a gain, although the statistics display rather broad tails and fluctuations. Using the full statistics of entanglement gain and loss we propose a classical stochastic random walk model with partial resetting for the entanglement dynamics, that we show to reproduce the full QJ dynamics. More importantly, the entanglement gain and loss picture clarifies the mutual role of quantum jumps and non-Hermitian evolution in the scaling of entanglement entropy in the steady-state, with direct implications for the associated MIPT.
@@
We showcase our new method by studying three models of monitored free-fermions: the Ising chain, the monitored SSH-chain and a model of free-fermions with $U(1)$ symmetry. These models are interesting for two reasons: first, they differ in their global symmetry, a fact which is believed to play a critical role in the stability of the MIPT; second, their entanglement transition in the no-click limit has been studied in detail Turkeshi and Schiró (2023); Gal et al. (2023). By solving the full stochastic QJ dynamics for these models and computing the entanglement entropy, we show that QJs have remarkably different impacts on their phase diagrams, as compared to the no-click limit. We understand this difference through a detailed analysis of the entanglement gain and loss, which reveals the crucial role of quantum jumps in renormalizing the effective non-Hermitian dynamics with respect to the bare no-click limit. We show that this renormalization, or its lack thereof, can naturally explain the different behaviors of quantum jumps in the three models and provide a criterion for the relevance of the no-click limit. Finally, we demonstrate how the the entanglement gain/loss picture and the associated classical model remain valid even in presence of interactions breaking gaussianity.
@@
The paper is structured as follows. In Sec. II we review the background material: we introduce the QJ measurement protocol and the models we will considered throughout this work as well as the quantities we will use to characterize them. In Sec. III we present the main quantity of interest for this work, namely the statistics of entanglement gain and loss, that we use to construct a classical stochastic model for the entanglement dynamics. In Sec. IV we present several applications of this tool to the study of MIPT in models of monitored free fermions. In particular we use the insights from this model to discuss the role of jumps and non-Hermitian dynamics on the scaling of the entanglement and the MIPT. Finally Sec. VI contains our conclusions and future perspectives. In the Appendixes we provide further methodological details and results relevant for our work.
Figure 1: Panel (a-b): cartoon of the setup. A monitored fermionic chain evolving under quantum jump protocol, characterized by a deterministic evolution driven by a non-Hermitian Hamiltonian $H_{\rm eff}$ (a) and stochastic Quantum Jumps (b). Panel (c): We are interested in the stochastic dynamics of the wave function and in particular in its entanglement structure, as measured from the bipartite entanglement entropy $S_{\xi}(t)$ (see panel a for the partition). Panel (d): typical quantum trajectory evolutions (10 realisations) for the entanglement entropy and its average over the measurement noise $\xi$.
II Background
@@
In this Section we summarize relevant background material, including the Quantum Jump protocol to continuously monitor a quantum system and its sampling via Montecarlo Wave Function method and the Waiting Time-Distribution. We then define the models we focus on throughout the paper and the main quantities we use to characterize their dynamics.
II.1 Monitoring by Quantum Jumps
In this work we are interested in the dynamics of continuously monitored quantum systems Wiseman and Milburn (2009). The set up we have in mind is sketched in Fig. (1), namely a fermionic chain where each lattice site is coupled to a measurement apparatus that weakly and continuously monitor some local observable of the chain. We are interested in the dynamics of the system wavefunction conditioned to the measurement outcomes, which realizes a so called quantum trajectory. Different measurement protocols give rise to different types of stochastic evolutions. Here we focus on the Quantum Jump (QJ) protocol Ueda (1990); Dalibard et al. (1992); Gardiner et al. (1992), corresponding to the experimentally relevant photo-counting process, and which is described by the following stochastic Schrodinger equation
$d|\Psi(t)\rangle$ $=-idt\left\{H-\frac{i}{2}\sum_{i}(L_{i}^{\dagger}L_{i}-\langle L_{i}^{\dagger}L_{i}\rangle_{t})\right\}|\Psi(t)\rangle$
$\ +\sum_{i}d\xi_{i}\left\{\frac{L_{i}}{\sqrt{\langle L_{i}^{\dagger}L_{i}\rangle}}-1\right\}|\Psi(t)\rangle.$ (1)
where $H$ is the Hamiltonian of the fermionic chain (we will give explicit expression in Sec. II.3), $L_{i},L^{\dagger}_{i}$ are the jump operators defined on each lattice site which describe the measurement process and $d\xi_{i}\in\{0,1\}$ is an increment for the in-homogeneous Poisson process with average ${P(d\xi_{i}=1)=dt\langle\psi(t)|L_{i}^{\dagger}L_{i}|\psi(t)\rangle}.$
@@ II.1 Monitoring by Quantum Jumps
The dynamics in Eq. (II.1) is composed of a deterministic part (for $d\xi_{i}=0$) which corresponds to an effective non-Hermitian Hamiltonian $H_{\rm eff}$ given by
$H_{\mathrm{eff}}=H-\frac{i}{2}\sum_{i}L_{i}^{\dagger}L_{i}$ (2)
@@
and a stochastic one, the last term in Eq. (II.1), due to the action of QJs. The effective Hamiltonian $H_{\rm eff}$ is non-Hermitian because of the measurement back-action, encoded in the last term of Eq. (2). Its role is to control the dynamics in the so called no-click limit, when no jumps happen during the quantum trajectory, as well as the evolution in-between two subsequent quantum jumps. We note that the evolution of the system is state-dependent (thus nonlinear); see the counter-term appearing in Eq. (II.1), to ensure the normalization.
The QJ evolution (II.1) is solved using Monte Carlo methods Daley (2014), either via a first-order integration scheme which introduces an explicit discretization $dt$ to sample the QJs or using higher-order schemes. The former approach has the drawback of not having a natural way to control the accuracy of the simulation, which is instead empirically benchmarked by considering different choices of $dt$ and checking that there is not qualitative difference in disorder averages. We overcome this limitation considering a higher-order scheme that is based on sampling the time at which subsequent jumps happen Breuer and Petruccione (2002); Daley (2014); Radaelli et al. (2023), i.e., using the cumulative waiting-time distribution defined as
@@
Specifically, we proceed by iterating the following loop : (i) Assume a normalised state $|\Psi(t_{i})\rangle$ is reached at time $t_{i}$ (possibly the initial time of the dynamics, in which case $|\Psi(t_{i})\rangle$ is the initial state). (ii) Extract a random number $r$ uniformly distributed in $[0,1]$ and find the random waiting time $\tau$ for the next jump by solving the equation $r=1-F[\Psi(t),\tau]$.
@@
(iv) At time $t+\tau$ a quantum jump occurs. As before, the output channel is chosen splitting the [0,1] interval into segments of size $\langle\Psi(t+\tau)|L^{\dagger}_{i}L_{i}|\Psi(t+\tau)\rangle$ and checking in which one a uniformly drawn random number falls. The jump is immediate and the post-measurement state after a jump at site $j$ reads
$|\Psi(t+\tau^{+})\rangle=\frac{L_{j}|\Psi(t+\tau)\rangle}{\sqrt{\langle\Psi(t+\tau)|L_{j}^{\dagger}L_{j}|\Psi(t+\tau)\rangle}}.$ (5)
@@
In a nutshell, a quantum trajectory is specified as a sequence of non-Hermitian quantum quenches interspersed with discontinuous jumps (see Fig. 1 (c)), that set the initial conditions for the forecoming integration steps.
An advantage of this approach is to access directly the waiting-time distribution (WTD) or delay function Cohen-Tannoudji and Dalibard (1986); Daley (2014); Landi et al. (2023), which encodes the probability distribution of the times between QJs. Its behavior will be discussed for the the specific model of interest later on in the manuscript.
II.2 Average vs Conditional Dynamics and Entanglement Entropy
The stochastic Schrödinger equation (II.1) describes the evolution of the conditional state, also called a quantum trajectory. By averaging the density matrix $\rho_{\xi}(t)=|\psi_{\xi}(t)\rangle\langle\psi_{\xi}(t)|$ over the stochastic noise $\xi$ in Eq. (II.1) one recovers a Lindblad master equation for the average state (This time dependent random variable $\mathbf{\xi}$ is used to label the different trajectories). These two descriptions are equivalent for what concerns simple observables, which are linear functionals of the conditional state, and indeed quantum trajectories are also called unravellings of the Lindblad master equation Carmichael (1999). The stochastic dynamics however contains a richer information: if one is interested in quantities which are non-linear in the state and sensitive to higher moments of the density matrix, then averaging over the measurement noise gives rise to physics which is not captured by the Lindblad/average state. A relevant example is provided by the (von Neumann) entanglement entropy, defined as  Calabrese and Cardy (2004); Amico et al. (2008)
$S_{\mathbf{\xi}}(t)=-\mathrm{tr}_{A}\left[\rho_{\mathbf{\xi}}^{A}(t)\ln\rho_{\mathbf{\xi}}^{A}(t)\right]\;,$ (6)